An infinite charges each of charge q are placed along x-axis at x=1,x=2,x=4,x=8,...Find the potential at x=0 due to this set of charges.What will be the potential if the consequent charges have opposite side?

Grade:11

1 Answers

Vikas TU
7 years ago
Potential at origin would be:
Vnet = Kq(1/1 + ½ + 1/4+ 1/8 + ….....+ infinity)
= Kq((1/(1-1/2)) = > 2Kq
For opposite side the potential would remain same as the distance would be constant and hence the potential is the scalar quantity.